# Cognitive Foundations for Reasoning and Their Manifestation in LLMs

### v1

@@ 3.4 Comparison with Humans
Distribution of Elements.
To complement our analysis of cognitive element presence and structure across models and problem types, we compare 30 manually annotated reasoning traces from both humans and LLMs. Figure 7 presents elements annotated as strongly present, excluding partial occurrences. Human traces exhibit significantly higher rates of abstract cognitive elements, such as self-awareness ($49\%$ vs. $19\%$), abstraction ($54\%$ vs. $36\%$). Notably, reasoning representations were frequently marked as partially present across both groups, particularly for humans, likely reflecting the more implicit nature of their internal knowledge representations and reasoning processes. Conversely, LLMs demonstrate greater reliance on backward chaining (navigating a problem by reasoning from goals to prerequisites; less common in human problem-solving), and higher productivity scores (potentially attributable to LLMs’ tendency to externalize intermediate reasoning steps that humans often leave implicit).
@@
Table 3 illustrates the difference between human and LLM reasoning traces. As shown, Humans are quicker to invoke conceptual processing and abstraction in logical problem solving, leading to significantly shorter reasoning traces. LLM’s on the other hand, often resort to surface level reiteration and enumeration. While LLMs try to reason abstractly and verify their reasoning, our Human annotators have repeatedly observed that they fail at learning from previous verifications in subsequent problem solving attempts, and they often repeat verification and backtracking on claims and paths that have already been explored. On more open ended questions, or questions that require complex factual recall, humans tend to invoke higher order behaviors such as self-awareness and strategy selection, leading to longer traces then LLMs, which seem to rely more on factual recall of relevant infromation.
@@ 4 Eliciting Cognitive Reasoning Structures
Section 3.3.2 established a methodology for extracting consensus subgraphs that encode the optimal hierarchical and temporal sequencing of reasoning behaviors for each problem type. We now investigate whether these empirically-derived cognitive structures can provide effective test-time reasoning guidance, steering models toward statistically successful reasoning patterns and thereby improving task performance.
@@
Table 4 reveals substantial heterogeneity in how models respond to cognitive structure guidance. Modern, capable reasoning models (particularly the Qwen3 family and larger R1-Distill variants) demonstrate significant improvements, with gains reaching up to 60% on ill-structured problem types (e.g., +60.0% for Qwen3-14B on dilemmas, +60.0% for R1-Distill-Qwen-32B on dilemmas). Notably, these improvements concentrate most strongly on complex, open-ended problems (dilemmas, case analysis, diagnostic reasoning) where explicit structural scaffolding appears most beneficial.
However, this effectiveness is highly dependent on model capacity and architectural sophistication. Smaller or less capable models, particularly Hermes-3-Llama-3-8B and DeepScaleR-1.5B, show pronounced performance degradation across most problem types, with losses exceeding 50% in several categories (e.g., -72.0% for both models on algorithmic problems). This suggests a capability threshold: models must possess sufficient reasoning flexibility and instruction-following ability to productively adapt their processes to detailed structural guidance. Below this threshold, explicit scaffolding appears to constrain rather than enhance reasoning, potentially by overwhelming limited cognitive resources or conflicting with ingrained problem-solving heuristics.
The pattern of improvements also illuminates problem-type specificity. Well-structured problems (algorithmic, rule-using) show more modest or even negative effects across models, while ill-structured problems exhibit the strongest positive responses in capable systems. This asymmetry likely reflects two factors: (1) well-structured problems may already elicit relatively optimal reasoning patterns from models, leaving less room for improvement, and (2) ill-structured problems benefit more from explicit organizational scaffolding that helps models navigate ambiguous problem spaces.
@@ 5 Cognitive Element Considerations in LLM Research Design
To contextualize the behavioral presence observed above in Section 3.3.1, we examine how contemporary LLM research conceptualizes "reasoning" as a design choice. With growing interest in the development and analysis of models with "diverse and strong reasoning" abilities, we aimed to understand what cognitive elements are currently supported and which remain underexplored. For this analysis, we scraped arXiv [^3] https://arxiv.org/ papers using keyword-based queries applied to titles and abstracts. We searched using general reasoning-related keywords (e.g., "large language model reasoning," "LLM cognitive behaviors," "LLM thinking") and for each capability in our taxonomy we added an additional query of the form "large language model <capability>", retrieving the top ten papers for each general query (in order to maximize precision) and 100 papers for each behavior-specific query.
@@
We then annotated this set of papers using both two human annotators and GPT-4.1 (all three achieve moderate agreement $ICC_{3k}=0.593$; we iteratively improve annotation prompts based on human feedback). For each paper–capability pair, we record whether a behavior is explicitly targeted (via evaluation objectives or architecture), implicitly encouraged (dataset structure, demonstration style), or not incorporated.
@@
Results from this analysis displayed in Figure 2 reveal substantial imbalance in the types of reasoning behaviors emphasized across LLM reasoning research. The dominant cognitive elements (context awareness: 70% of papers, decomposition and integration: 60%, knowledge structure alignment: 56%) align with linear procedural step-by-step reasoning that is straightforward to evaluate. Sequential organization (54%), pattern recognition (51%), and abstraction (46%) similarly emphasize forward-moving compositional patterns. In contrast, behaviors enabling flexible problem-solving receive far less attention: self-awareness appears in only 16% of papers, spatial organization in 10%, and temporal organization in 22%.
@@
Comparing paper-level design intentions to observed model behaviors reveals three key discrepancies. First, design-behavior gaps: Compositionality appears in 38% of papers yet shows inconsistent presence in reasoning traces. Context alignment and knowledge structure alignment are frequently targeted (47% and 56%) but models struggle to maintain these consistently across problem types. Second, emergent but under-theorized behaviors: Profundity appears in only 16% of papers but manifests consistently in model outputs, suggesting models develop meta-cognitive patterns not explicitly designed for. Third, systematically neglected capabilities: Self-awareness (16%), temporal organization (22%), ordinal organization (27%), and spatial organization (10%) represent sophisticated cognitive structures essential for non-linear thinking, mental simulation, and metacognitive monitoring. These capabilities receive minimal research attention and fail to emerge spontaneously in model behavior.
@@
This synthesis reveals a critical bottleneck. Current LLM reasoning research operates within a narrow conceptual vocabulary privileging linear, compositional behaviors amenable to straightforward evaluation. Expanding theoretical foundations beyond step-by-step decomposition toward richer cognitive taxonomies may be essential for developing more sophisticated reasoning capabilities.
@@ 6 Opportunities and Challenges
Our proposed cognitive foundations taxonomy enables systematic characterization of reasoning processes, allowing us to answer questions like which cognitive elements appear in models?, how behaviors unfold during reasoning? and which behavioral patterns correlate with success? Overall, our analyses expose fundamental gaps: we cannot know which training produces which cognitive capabilities a priori, cannot ensure behaviors transfer beyond training distributions, and cannot validate whether observed patterns reflect genuine cognitive mechanisms or spurious reasoning shortcuts. The cognitively inspired test-time reasoning guidance demonstrates that this understanding is actionable, in which successful behavioral patterns can be elicited to improve performance on ill-structured problems (Table 4). However, even in this case, it remains unclear whether our guidance enables genuine deployment of latent capabilities or simply helps models retrieve cached reasoning patterns from training data.
@@
Current training paradigms lack predictive theories connecting procedures to emergent reasoning capabilities. Meta-cognitive monitoring correlates strongly with success on ill-structured problems (Figure 4), yet appears in only 8% of reasoning traces and 8% of papers (Figure 2). Our comparison across 16 models shows dramatic variation in response to cognitive scaffolding (Table 4), but we cannot predict these differences from architecture or training details. Post-hoc analyses reveal that RL induces verification (gandhi2025cognitive; snell2024scaling) but not representational restructuring or meta-cognitive monitoring, while process supervision (lightman2023lets; uesato2022solving) and chain-of-thought prompting (wei2022chain; kojima2022large; wang2023selfconsistency; yao2023tree) elicit latent but not spontaneous behaviors. We cannot explain why RL produces verification but not meta-cognitive monitoring, or whether architectural prerequisites exist for specific behaviors versus emerging from scale (le2025reasoning; das2025can).
@@
These principles enable predictive theories of capability emergence. If procedural behaviors emerge from repetition while meta-cognitive behaviors require reflection, we can predict which training paradigms will produce which capabilities before running experiments. If transferable reasoning requires structural variation across surface features, we can predict that training on homogeneous problem distributions will fail to produce robust generalization. Our taxonomy provides measurement infrastructure for testing these predictions; cognitive theories provide the explanatory framework linking training characteristics to behavioral outcomes. This enables shifting from post-hoc observation to theory-driven experimentation.
@@
Our framework identifies observable behavioral patterns, but the same surface behavior can arise from fundamentally different underlying processes (chomsky2014aspects). Our human-LLM comparison illustrates this: both reach correct answers, yet humans employ hierarchical nesting and meta-cognitive loops while models use shallow sequential chaining (Figure LABEL:fig:transition_graph). Recent work shows models produce correct reasoning chains while internally representing different processes (turpin2023language) and that chain-of-thought may be post-hoc rationalization (lanham2023measuring). These alternatives produce similar outputs but fundamentally different robustness and generalization.
@@
Developing validation frameworks that test these signatures requires moving beyond "does behavior X appear?" to "does X transfer systematically, remain robust to perturbation, and compose flexibly?" This demands systematic probing studies measuring transfer and robustness (belinkov2022probing), causal interventions manipulating specific capabilities while measuring downstream effects, and mechanistic interpretability connecting internal representations to behavioral function (elhage2021mathematical; nanda2023progress). Our taxonomy identifies which behaviors to validate; cognitive theories specify what properties distinguish genuine capabilities from spurious shortcuts.
@@
Models serve as computational implementations of cognitive hypotheses that can be systematically manipulated and tested (griffiths2019doing; saxe2021if). We can examine how reasoning behaviors emerge during training (chang2024language; wei2022emergent), ablate architectural components to identify their functional roles (sternberg2009cognitive), and test theories of hierarchical planning and meta-cognition through interventions impossible with human subjects. Recent work uses models to test theories of semantic cognition (grand2022semantic), conceptual knowledge (patel2022bidirectional), and pragmatic reasoning (tessler2016pragmatic) at scales unachievable in traditional experiments.
@@ A.2.1 Extension of Jonassen’s Taxonomy
  • Factual Recall: Retrieving stored knowledge without requiring reasoning or problem-solving (e.g., "What is photosynthesis?" or "List the causes of WWI")
  • Creative/Expressive: Generating novel content judged by originality or aesthetic quality rather than convergence to a predetermined solution (e.g., "Write a poem" or "Draw how you feel")
@@ A.2.3 Classification Methodology
Each problem is classified through majority voting across three frontier language models (GPT-4o-mini, Gemini-2.5-Pro, and Claude-Sonnet-4.5). Each model independently classifies the problem using detailed annotation guidelines based on Jonassen’s (2000) structural criteria: goal clarity, solution determinacy, and domain constraints. Three-way disagreements occur in under 3% of cases and are adjudicated manually using these structural criteria. The complete classification prompt and guidelines are available in our code repository at https://github.com/stellalisy/CognitiveFoundations.
@@ A.2.4 Key Distinctions
This taxonomy enables systematic analysis of how problem structure affects behavioral manifestation and reasoning success across our dataset of 171,485 traces.
@@ A.3 Accuracy Analysis
Table 5 presents the complete accuracy results for all 14 text reasoning models across the 13 problem types in our taxonomy. Numbers in parentheses indicate the number of problems evaluated for each model-type pair. The models span five major architectural families: Qwen3 (Alibaba’s native thinking mode integration), DeepSeek-R1 and its distillations (knowledge transfer from 671B teacher), OpenThinker (data-quality focused), DeepScaleR (efficient RL), s1.1 (Qwen-based efficient training), and DeepHermes-3 (hybrid reasoning with user-controlled depth).

### v2

@@ 3.4 Comparison with Humans
Distribution of Elements.
To complement our analysis of cognitive element presence and structure across models and problem types, we compare 30 manually annotated reasoning traces from both humans and LLMs. Note that our human participants are educated adults; comparing LLMs to human developmental trajectories (children acquiring reasoning skills) may provide different insights, as cognitive elements emerge progressively through development (goswami1996analogical; sandberg2011development). Figure 7 presents elements annotated as strongly
@@
present or partially present. Human traces exhibit significantly higher strongly present rates of abstract cognitive elements, such as heightself-awareness ($49\%$ vs. $19\%$) and heightabstraction ($54\%$ vs. $36\%$). Notably, heightreasoning representations were frequently marked as partially present across both groups, particularly for humans, likely reflecting the more implicit nature of their internal knowledge representations and reasoning processes. Conversely, LLMs demonstrate greater reliance on heightbackward chaining (navigating a problem by reasoning from goals to prerequisites; less common in human problem-solving), and heightproductivity which potentially attributable to LLMs’ tendency to externalize intermediate reasoning steps that humans often leave implicit.
@@
Table 3 illustrates the difference between human and LLM reasoning traces. As shown, Humans are quicker to invoke heightconceptual processing and heightabstraction in logical problem-solving, leading to significantly shorter reasoning traces. LLM’s on the other hand, often resort to surface level reiteration and enumeration. While LLMs try to reason abstractly and verify their reasoning, manual annotations have repeatedly observed that they fail at learning from previous verifications in subsequent problem-solving attempts, and they often repeat heightverification and heightbacktracking on claims and paths that have already been explored. On more open-ended questions, or questions that require complex factual recall, humans tend to invoke higher order behaviors such as heightself-awareness and heightstrategy selection, leading to longer traces than LLMs, which seem to rely more on factual recall of relevant information.
@@ 4 Eliciting Cognitive Reasoning Structures
Section 3.3.2 established a methodology for extracting reasoning structures that encode the successful hierarchical and temporal sequencing of behavioral manifestations of cognitive elements for each problem type. We now apply these empirically-derived cognitive structures to provide effective test-time reasoning guidance, steering models toward successful reasoning patterns and thereby improving task performance.
@@
Table 4 reveals substantial heterogeneity in how models respond to cognitive structure guidance. Modern, capable reasoning models (particularly the Qwen3 family and larger R1-Distill variants) demonstrate significant improvements, with gains reaching up to 66.7% on ill-structured problem types (e.g., +66.7% for Qwen3-7B on diagnosis, +60.0% for Qwen3-14B on dilemmas, +60.0% for R1-Distill-Qwen-32B on dilemmas). Notably, these improvements concentrate most strongly on complex, open-ended problems (dilemmas, case analysis, diagnosis) where explicit structural scaffolding appears most beneficial.
However, this effectiveness is highly dependent on model capacity and architectural sophistication. Smaller or less capable models, particularly Hermes-3-Llama-3-8B and DeepScaleR-1.5B, show pronounced performance degradation across most problem types, with losses exceeding 50% in several categories (e.g., -72.0% for both on algorithmic problems). This suggests a capability threshold: models must possess sufficient reasoning flexibility and instruction-following ability to productively adapt their processes to detailed structural guidance. Below this threshold, explicit scaffolding appears to constrain rather than enhance reasoning, potentially by overwhelming limited cognitive resources or conflicting with ingrained problem-solving heuristics.
The pattern of improvements also illuminates problem-type specificity. Well-structured problems (algorithmic, rule-using) show more modest or even negative effects across models, while ill-structured problems exhibit the strongest positive responses in capable systems. This directly echoes the observation from Figure 4, where the well-structured problems show lower behavioral presence and the cognitive elements are less predictive of task success. On the other hand, the ill-structured problems benefit more from explicit organizational scaffolding that helps models navigate ambiguous problem spaces.
@@ 5 Cognitive Element Considerations in LLM Research Design
To contextualize the behavioral presence observed above in Section 3.3.1, we examine how contemporary LLM research conceptualizes “reasoning” as a design choice. With growing interest in the development and analysis of models with “diverse and strong reasoning” abilities, we aim to understand what cognitive elements are currently supported and which remain underexplored. For this analysis, we scrape arXiv [^3] https://arxiv.org/ papers using keyword-based queries applied to titles and abstracts. We search using general reasoning-related keywords (e.g., “LLM reasoning,” “LLM cognitive behaviors,” “LLM thinking”) and for each capability in our taxonomy we add an additional query of the form “LLM <capability>,” retrieving the top ten papers for each general query (in order to maximize precision) and 100 papers for each behavior-specific query.
@@
We annotate a subset of papers using two human annotators and GPT-4.1 (all three achieve moderate agreement $ICC_{3k}=0.593$; we iteratively improve annotation prompts based on human feedback), and then the entire set of 1598 papers using GPT-4.1 with the iterated prompts. For each paper–capability pair, we record whether a cognitive element is explicitly targeted (via evaluation objectives or architecture), implicitly encouraged (dataset structure, demonstration style), or not incorporated.
@@
Results from this analysis displayed in Figure 2 reveal substantial imbalance in the types of cognitive elements emphasized across LLM reasoning research. The dominant cognitive elements (context awareness: 70% of papers, decomposition and integration: 60%, knowledge structure alignment: 56%) align with linear procedural step-by-step reasoning that is straightforward to evaluate. Sequential organization (54%), pattern recognition (51%), and abstraction (46%) similarly emphasize forward-moving compositional patterns. In contrast, elements enabling flexible problem-solving receive far less attention: self-awareness appears in only 16% of papers, spatial organization in 10%, and temporal organization in 22%.
@@
Comparing paper-level design intentions to observed model behaviors reveals three key discrepancies. First, design-behavior gaps: Compositionality appears in 38% of papers yet shows inconsistent presence in reasoning traces. Context alignment and knowledge structure alignment are frequently targeted (47% and 56%) but models struggle to maintain these consistently across problem types. Second, emergent but under-theorized behaviors: Profundity appears in only 16% of papers, but manifests consistently in model outputs, suggesting models develop meta-cognitive patterns not explicitly designed for. Third, systematically neglected capabilities: Self-awareness (16%), temporal organization (22%), ordinal organization (27%), and spatial organization (10%) represent sophisticated cognitive structures essential for non-linear thinking, mental simulation, and metacognitive monitoring. These capabilities receive minimal research attention and fail to emerge spontaneously in model behavior.
@@
This synthesis reveals a critical bottleneck. Current LLM reasoning research operates within a narrow conceptual vocabulary, privileging linear, compositional behaviors amenable to straightforward evaluation. Expanding theoretical foundations beyond step-by-step decomposition toward richer cognitive taxonomies may be essential for developing more sophisticated reasoning capabilities.
@@ 6 Opportunities and Challenges
Our proposed cognitive foundations taxonomy enables systematic characterization of reasoning processes, allowing us to answer questions like which cognitive elements appear in models?, how cognitive elements unfold during reasoning? and which behavioral patterns correlate with success? Overall, our analyses expose fundamental gaps: we cannot know which training produces which cognitive capabilities a priori, cannot ensure cognitive elements transfer beyond training distributions, and cannot validate whether observed patterns reflect genuine cognitive mechanisms or spurious reasoning shortcuts. The cognitively inspired test-time reasoning guidance demonstrates that this understanding is actionable, in which successful behavioral patterns can be elicited to improve performance on ill-structured problems (Table 4). However, even in this case, it remains unclear whether our guidance enables genuine deployment of latent capabilities or simply helps models retrieve cached reasoning patterns from training data.
@@
Current training paradigms lack predictive theories connecting procedures to emergent reasoning capabilities. Meta-cognitive monitoring correlates strongly with success on ill-structured problems (Figure 4), yet appears in only 8% of reasoning traces and 8% of papers (Figure 2). Our comparison across 16 models shows dramatic variation in response to cognitive scaffolding (Table 4), but we cannot predict these differences from architecture or training details. Post-hoc analyses reveal that RL induces verification (gandhi2025cognitive; snell2024scaling) but not representational restructuring or meta-cognitive monitoring, while process supervision (lightman2023lets; uesato2022solving) and chain-of-thought prompting (wei2022chain; kojima2022large; wang2023selfconsistency; yao2023tree) elicit latent but not spontaneous behaviors. Our framework enables testing whether specific cognitive elements require architectural prerequisites or emerge from scale (le2025reasoning; das2025can), and which training procedures produce which cognitive profiles. These patterns may reflect how autoregressive next-token prediction biases models toward sequential processing (bachmann2024pitfalls; alberghi2025bias), or how outcome-based rewards prioritize answer correctness over reasoning soundness (ye2025beyond). However, our analysis examines only transformer-based language models; attributing causes requires systematic comparison across architectures and training paradigms.
@@
These principles enable predictive theories of capability emergence. If procedural cognitive elements emerge from repetition while meta-cognitive elements require reflection, we can predict which training paradigms will produce which capabilities before running experiments. If transferable reasoning requires structural variation across surface features, we can predict that training on homogeneous problem distributions will fail to produce robust generalization. Our taxonomy provides measurement infrastructure for testing these predictions; cognitive theories provide the explanatory framework linking training characteristics to behavioral outcomes. This enables shifting from post-hoc observation to theory-driven experimentation.
@@
Our framework identifies observable behavioral patterns, but the same surface behavior can arise from fundamentally different underlying processes (chomsky2014aspects). Our human-LLM comparison illustrates this: both reach correct answers, yet humans employ hierarchical nesting and meta-cognitive loops while models use shallow sequential chaining (Figure 6). Recent work shows models produce correct reasoning chains while internally representing different processes (turpin2023language) and that chain-of-thought may be post-hoc rationalization (lanham2023measuring). These alternatives produce similar outputs but fundamentally different robustness and generalization.
@@
Developing validation frameworks that test these signatures requires moving beyond “does behavior X appear?” to “does X transfer systematically, remain robust to perturbation, and compose flexibly?” This demands systematic probing studies measuring transfer and robustness (belinkov2022probing), causal interventions manipulating specific capabilities while measuring downstream effects, and mechanistic interpretability connecting internal representations to behavioral function (elhage2021mathematical; nanda2023progress). Our taxonomy identifies which behaviors to validate; cognitive theories specify what properties distinguish genuine capabilities from spurious shortcuts.
@@
Models serve as computational implementations of cognitive hypotheses that can be systematically manipulated and tested (griffiths2019doing; saxe2021if). We can examine how cognitvie elements emerge during training (chang2024language; wei2022emergent), ablate architectural components to identify their functional roles (sternberg2009cognitive), and test theories of hierarchical planning and meta-cognition through interventions impossible with human subjects. Recent work uses models to test theories of semantic cognition (grand2022semantic), conceptual knowledge (patel2022bidirectional), and pragmatic reasoning (tessler2016pragmatic) at scales unachievable in traditional experiments.
@@ A.2.1 Extension of Jonassen’s Taxonomy
  • Factual Recall: Retrieving stored knowledge without requiring reasoning or problem-solving (e.g., “What is photosynthesis?” or “List the causes of WWI”)
  • Creative/Expressive: Generating novel content judged by originality or aesthetic quality rather than convergence to a predetermined solution (e.g., “Write a poem” or “Draw how you feel”)
@@ A.2.3 Classification Methodology
Each problem is classified through majority voting across three frontier LLMs (GPT-4o-mini, Gemini-2.5-Pro, and Claude-Sonnet-4.5). Each model independently classifies the problem using detailed annotation guidelines based on Jonassen’s (2000) structural criteria: goal clarity, solution determinacy, and domain constraints. Three-way disagreements occur in under 3% of cases and are adjudicated manually using these structural criteria. The complete classification prompt and guidelines are available in our code repository at https://github.com/stellalisy/CognitiveFoundations.
@@ A.2.4 Key Distinctions
This taxonomy enables systematic analysis of how problem structure affects behavioral manifestation and reasoning success across our dataset of 192,709 traces.
@@ A.3 Accuracy Analysis
Table 5 presents the complete accuracy results for all 16 text reasoning models across the 13 problem types in our taxonomy. Numbers in parentheses indicate the number of problems evaluated for each model-type pair. The models span five major architectural families: Qwen3 (Alibaba’s native thinking mode integration), DeepSeek-R1 and its distillations (knowledge transfer from 671B teacher), OpenThinker (data-quality focused), DeepScaleR (efficient RL), s1.1 (Qwen-based efficient training), and DeepHermes-3 (hybrid reasoning with user-controlled depth).